Milk that is 4% butterfat is mixed with milk that is 1% butterfat to obtain 18 gallons of milk that is 2% butterfat. How many gallons of each type of milk are needed?
I don't even know where to start, PLEASE HELP!!!!!

Answer by Paul(988)

4x+1(18-x)=18(2)
4x+18-x=36
3x=18
x=6

18-6=12

Hence about 6 gallons of 4% is needed, and about 12 gallons of 1% is needed.
Paul.